i need to graph x+3y=9
Answer by Alan3354(69443)

i need to graph x+3y=9
---------------
It's a straight line, so you need 2 points.
x = 0 --> y = 3
(0,3) point 1
----------------
y = 0, x = 9
(9,0) point 2
Draw a line thru the 2 points.